OpenStudy (anonymous):

HELP! I GIVE MEDALS!!! Which of the following represents 3x - 5y + 10 = 0 written in slope-intercept form? y = -3/5x + 2 y = 3/5x - 2 y = 3/5x + 2

OpenStudy (bibby):

isolate the y variable -5y = -3x-10 How would you proceed?

OpenStudy (anonymous):

ummmm i have no idea

OpenStudy (bibby):

You're solving for y. How do you get y alone?

OpenStudy (anonymous):

divide it by itself?

OpenStudy (bibby):

eh close enough. divide by -5 to isolate y.\[y = \frac{ -3x-10 }{ -5 } = \frac{ -3 }{ -5 }x - \frac{ 10 }{ -5 }= \frac{ -3 }{ -5 }x - \frac{ 10 }{ -5 }\]

OpenStudy (shamil98):

Slope-intercept form is: \[\Huge y = mx + b\] Your equation is \[\huge 3x - 5y + 10 = 0\] You need to solve the equation for y to get that form.

OpenStudy (shamil98):

As bibby has shown, but you can simplify it further.. do that to get your final answer..

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

\[y=3/5x+2\]

OpenStudy (bibby):

my dad is a mod I reported you

OpenStudy (shamil98):

-(10/-5) = what?

OpenStudy (bibby):

this is my gay dad

OpenStudy (anonymous):

....

OpenStudy (bibby):

- -2 = +2

OpenStudy (shamil98):

Fed dat answer eh